A basket of fruits contains eight (8) apples and ten (10) oranges. Half of the apples and half of the oranges are rotten. If one (1) fruit is chosen at random, what is the probability that a rotten apple or an orange is chosen? 


1
Expert's answer
2022-06-01T11:07:30-0400
"P(rotten\ apple\ or\ rotten\ orange)=\dfrac{8/2+10/2}{8+10}""=\dfrac{1}{2}""P(rotten\ apple\ or\ any\ orange)=\dfrac{8/2+10}{8+10}""=\dfrac{7}{9}"